in triangle EFG, FH is the bisector of angle F. if EH=6 and HG= 9, find the ratio of FG:EF.
Answer by venugopalramana(3286)

IN A TRIANGLE THE ANGLE BISECTOR OF ANY ANGLE DIVIDES THE OPPOSITE SIDE IN THE RATIO OF 2 ADJACENT SIDES OF THE TRIANGLE
HERE EFG IS A TRIANGLE, FH is the bisector of angle F.IT MEETS THE OPPOSITE SIDE EG AT H.HENCE AS PER THE THEOREM WE HAVE
EH/HG=FE/FG
if EH=6 and HG= 9,
HENCE EH/HG=6/9=FE/FG
FG/FE=9/6=3/2
the ratio of FG:EF=3:2